OpenStudy (anonymous):

Can someone check my work and see if I did it wrong or right? If wrong please explain to me the right answer thanks! Medal+Fan (:

OpenStudy (anonymous):

@tHe_FiZiCx99 @ml1042903

OpenStudy (superdavesuper):

1st one is correct....2nd one i think u made a minor error in ur math; would u like to try again n see if u can figure it out? :)

OpenStudy (superdavesuper):

the general eqn for 2 points is (y-y1)/(x-x1)=(y1-y1)/(x1-x2) so (y-3)/(x--1)=(3-11)/(-1-1) simplify n re-arrange...

OpenStudy (anonymous):

A?

OpenStudy (superdavesuper):

Nope....check out my eqn above plz :)

OpenStudy (superdavesuper):

(y-3)/(x--1)=(3-11)/(-1-1) (y-3)/(x+1)=-8/-2....

OpenStudy (anonymous):

D? sorry I don't really get this stuff

OpenStudy (superdavesuper):

Bingo - good job :)

OpenStudy (the_fizicx99):

The first one is correct, you made a small error on the second. First find the slope, M = (y2-y1) / (X2-x1) M = 11-3 / 1 - (-1) M = 8 / 2 M = 4 So the slope is 4 Now write the equation, y-y1=m(x-x1) y-3=4(x-(-1)) y-3=4x + 4 +3 +3 y = 4x + 7 So rewrite it

OpenStudy (anonymous):

Thank you both! Oh god not sure who to give the medal too lol can you check two of my other questions?

OpenStudy (the_fizicx99):

sure

OpenStudy (superdavesuper):

i will let some1 else help u now @mysticghost :)

OpenStudy (anonymous):

Okay thanks for helping me before

OpenStudy (anonymous):

@tHe_FiZiCx99 posted it ^ just help me when you can thanks (:

OpenStudy (the_fizicx99):

K um, parrallel have the same slope, perpendicular is the reciprocal, so flip it and multiply it by -1 You have the slope, so for parallel, just set up the equation. y-y1=m(x-x1) y-9=1/4(x-(-8)) y-9=1/4x + 2 +9 +9 y = 1/4x + 11 <-- this is the slope intercept form for parallel. Can you find the perpendicular one?

OpenStudy (the_fizicx99):

You already know the answer by just looking at it, but you're more than welcome to try, I'll correct you if your wrong :PP dw

OpenStudy (anonymous):

Is it D?

OpenStudy (the_fizicx99):

How did you get D ?

OpenStudy (anonymous):

well its the only one with a slop intercept for the parallel with y = 1/4x + 11 unless its A.

OpenStudy (the_fizicx99):

o.o lol but but.. i gave you the parallel you didn't try the perp one U_U

OpenStudy (anonymous):

yes i know lol but if thats the parallel then D is the only answer with that one so thats why i think its D but i also think it could be A.

OpenStudy (the_fizicx99):

It's D U_U A is -11 >.<

OpenStudy (anonymous):

I said it was D...

OpenStudy (the_fizicx99):

^ Idk, i guess i wanted to help you understand how to do both >.<

OpenStudy (anonymous):

Oh well thank you!

OpenStudy (the_fizicx99):

yw :]

OpenStudy (anonymous):

(: mind checking one more?

OpenStudy (the_fizicx99):

nope :p

OpenStudy (the_fizicx99):

It's all real numbers greater than 0 and less than 13

OpenStudy (the_fizicx99):

24 + 2x < 50 -24 -24 2x < 26 /2 /2 x < 13

OpenStudy (anonymous):

Ohh I think I get it thanks!

OpenStudy (the_fizicx99):

yw